### v1

# Towards Real-time Adaptive Anisotropic Image-to-mesh Conversion for Vascular Flow Simulations

Abstract
Presented is a path towards a fast and robust adaptive anisotropic mesh generation method that is designed to help streamline the discretization of complex vascular geometries within the Computational Fluid Dynamics (CFD) modeling process. The proposed method combines multiple software tools into a single pipeline to provide the following: (1) image-to-mesh conversion which satisfies quality, fidelity, and smoothness requirements, (2) the generation of a boundary layer grid over the high fidelity surface, (3) a parallel adaptive anisotropic meshing procedure which satisfies real-time requirements, and (4) robustness, which is satisfied by the pipeline’s ability to process segmented images and CAD models. The proposed approach is tested with two brain aneurysm cases and is shown to satisfy all the aforementioned requirements. The next steps are to fully parallelize the remaining components of the pipeline to maximize potential performance and to test its integration within a CFD vascular flow simulation. Just as the parallel anisotropic adaptation procedure was tested within aerospace CFD simulations using CAD models, the method is expected to provide accurate results for CFD vascular flow simulations in real-time when executed on multicore cc-NUMA architectures.
1 Introduction.
Interventional radiology involves minimally invasive, catheter-based, deployment of several vascular diseases throughout the body. This field has seen an explosion of device technology over the past decade ranging from stent-grafts for aortic aneurysms to valve replacement for diseased cardiac valves. Neuro-interventional radiology, which deals with diseases of the cerebral vasculature, has expanded from the treatment of brain aneurysms to removal of blood clots in large vessel occlusion strokes to embolization of the middle meningeal artery for treatment of chronic subdural hematomas to venous sinus stenting for treatment of idiopathic intracranial hypertension and pulsatile tinnitus. As hemodynamics are an intrinsic factor of vascular disease, there has been a concomitant increase in the number of studies using Computational Fluid Dynamics (CFD) and/or fluid-structure interactions to better understand disease prognosis or treatment outcomes after deployment of the pertinent medical device. For example, the transport of cardiogenic emboli causing ischemic stroke can be evaluated using imaging scans of the entire vascular branching network in each patient [51]. The recent introduction of miniaturized optical coherence tomography allows for high-resolution imaging of medical device features as well as the vascular wall in any given patient [53], which in turn facilitates high-resolution computational studies of the pathology. CFD may be able to provide clarity in the device treatment of cerebral venous disease, which has irregular morphologies and complex flow patterns [62].
Brain aneurysms (such as that shown in fig. 1) are abnormal focal dilations of intracranial arteries which when left untreated may rupture, resulting in fatal outcomes for patients. Aneurysms have been treated by neuro-interventional devices for nearly three decades, and hence, numerous CFD studies have been conducted on this disease. High resolution CFD simulations provide the opportunity to study hemodynamics in arteries, which can assist clinicians in determining an aneurysm’s potential to rupture and possible methods of treatment (such as the design of neurovascular stents) [52, 13, 17, 12, 36, 46]. Brain aneurysms that form at vascular bifurcations may be subjected to flow instabilities that can only be captured by high-fidelity CFD [41]. Stents used to treat brain aneurysms have wire sizes that can be as small as 25-30 microns with aneurysm sizes ranging from millimeters to centimeters. CFD studies on the effect of such devices (including overlapping devices) on intraneurysmal hemodynamics involve generation of volumetric meshes with elements at multiple sizes [24]. Processing patient-specific medical imaging data of cerebral aneurysms involves the discretization of these complex bodies in the form of smoothed free-form anisotropic body-conforming meshes. Mesh generation methods utilized in commercially available products work well when processing patient-specific cerebral aneurysm geometries; however, these methods do not always satisfy the time constraints of high resolution CFD simulations.
Figure 1: A) an angiogram showing a large $\sim$25 mm aneurysm on the right internal carotid artery; B) vasculature segmented in SimVascular [3], lofted in Rhinoceros [2], and imported into SolidWorks [4]; C) flow diverter stent simulated in Matlab with VMTK [9] input; D) vasculature and stent bodies imported into ADINA [1] for mesh generation (also appears in [25]).
These types of medical simulations can be classified into two categories: interactive and predictive. Interactive simulations require real-time mesh generation which can be used to allow surgical residents to practice performing a procedure within a virtual environment. Real-time mesh generation is also required for image registration applications, as they provide guidance to surgeons during operations [21]. Predictive simulations predict and optimize the outcome of an intervention using patient-specific pre-operative image data. Although predictive simulations don’t necessarily require real-time mesh generation, both types of simulations are time-sensitive given their impact on a patient’s treatment and associated costs.
This paper presents a path towards a fast and robust adaptive anisotropic mesh generation method that is designed to help streamline the discretization of complex vascular geometries and Computer-Aided Design (CAD)-based stents [25, 40] within the CFD modeling process. Several requirements should be addressed with regards to medical image-to-mesh (I2M) conversion: real-time mesh generation, robustness, fidelity, quality, and smoothness. Robustness concerns the software’s ability to process different input types such as CAD models (for example, of stents) and patient-specific medical images. Utilizing both is essential when running a CFD vascular flow simulation meant to aid in determining potential methods of treatment [40, 26, 25]. Fidelity measures the degree to which the mesh surface aligns with an image boundary. Quality (determined by the shape and size of mesh elements) and smoothness affect the accuracy of solutions for CFD simulations [10, 32]. Each of these requirements are independently challenging problems. The proposed approach addresses the robustness, fidelity, quality, and (to a degree) smoothness requirements (providing a smoothness of $C^{0}$, as opposed to other available real-time I2M methods that do not address smoothness [31]). Given that the image-to-mesh conversion and boundary layer generation steps are sequential, the proposed approach only addresses the real-time requirement with its parallel adaptive anisotropic mesh generation. After this feasibility study, the next steps will be to fully parallelize all components of the proposed approach in order to maximize potential performance.
The proposed method involves several steps: (1) utilize a method known as CBC3D [27] to discretize a segmented image using a Body-Centered Cubic lattice of high-quality tetrahedra and deform the generated mesh surfaces to their corresponding tissue boundaries to improve fidelity while maintaining quality, (2) construct an anisotropic metric tensor field, (3) generate a boundary layer grid outside the high-fidelity CBC3D mesh surface using a method known as Advancing Front Local Reconnection (AFLR) [48], and (4) utilize a multicore cc-NUMA-based mesh adaptation method known as CDT3D [64] to generate an anisotropic mesh volume using the metric tensor field and CBC3D isotropic mesh as a background mesh. We evaluate this approach using two aneurysm cases (a carotid cavernous aneurysm and middle cerebral artery bifurcation aneurysm) obtained from the Cerebrovascular Center for Research at Stony Brook University under Institutional Review Board approval. We show that the proposed approach meets the aforementioned I2M conversion requirements. Given that CDT3D also exhibited excellent performance and accurately captured features of underlying aerospace simulations when coupled with a solver (processing CAD data such as that shown in fig. 2), the end goal is to utilize this method within a vascular flow simulation.
Figure 2: A CDT3D mesh of an aerospace geometry generated from CAD data and its associated solution obtained from a CFD simulation are shown (also appears in [64]).
2 Related Work.
There are numerous methods that process medical data as input to create meshes. The type of meshes generated must be taken into account when considering the target application. Meshes can be classified into two categories - structured and unstructured. While some structured mesh generation methods have been shown to provide accurate results within CFD simulations [54, 38, 59], these types of methods can sometimes fail to capture features within complex geometries [11]. Unstructured meshes can be further classified into two categories - isotropic and anisotropic. Isotropic mesh generation methods are well suited for high curvature geometries. Anisotropic meshes contain high aspect ratio elements that include directional information (e.g., hemodynamics). Additionally, they may contain fewer elements in some cases compared to isotropic meshes generated from the same image. A mesh should ideally be as small as possible to reduce the memory and CPU requirements for the solver. These benefits make anisotropic meshes suitable for blood flow problems.
There are many methods that convert 3D images into 3D meshes. For a comprehensive overview of these methods, see [68]. Many 3D isotropic methods utilize Delaunay refinement techniques [14, 56, 34, 60, 16, 58, 18, 31, 63, 30]. A challenge with Delaunay refinement, which still remains an open problem, is that almost flat tetrahedra (slivers) can survive known heuristics designed to remove them [14, 18, 20]. These low quality elements introduce error when processed by a solver. Additional I2M methods utilize lattice space-tree (octree) decomposition techniques [15, 70, 69, 42, 19, 67, 71, 44, 37, 45, 57, 50] to provide meshes with adaptively-sized, high quality elements. Some also utilize the Dual Contouring technique to improve element quality [33, 71, 44, 37, 45]. Others warp the surfaces of the lattice to the boundaries of the object using either a mass-spring system, an FE constitutive model, or an optimization scheme [50, 46, 34]. Some methods focus on processing surface geometries of medical data to create structured Cartesian cut-cell meshes [55] and isotropic tetrahedral meshes [39]. [29] evaluates several of these sequential isotropic-based methods’ suitability for surgical simulations and non-rigid registration applications. There are methods which utilize parallel mesh generation techniques (with either CPUs [45, 54, 38, 59] or GPUs [45]). However, [54, 38, 59] all focus on generating Cartesian cut-cell meshes while [45] focuses on generating isotropic tetrahedral meshes, as with all the aforementioned Delaunay and lattice-based methods. [37] utilizes an anisotropic Giaquinta–Hildebrandt operator-based geometric flow method to smooth mesh surfaces while preserving surface and volume features. However, this smoothing is applied to isotropic meshes.
Some methods focus on converting medical images into anisotropic meshes. Although potentially useful for blood flow simulations, [22] focuses on using an anisotropic mesh adaptation procedure to construct a high-fidelity virtual model of a fractured pelvis, and then to build a physical replica (by means of additive manufacturing techniques) to represent the anatomical object from a segmented image. [16] generates high-quality adaptive hybrid meshes for fluid-solid geometries within cardiac environments. Following segmentation, surfaces between heart tissue and blood (and other non-cardiovascular tissue/fluids) are extracted and adapted. Anisotropic tetrahedra are generated for the interior of heart tissue, a prismatic boundary layer is generated for the interface between tissue and blood, and TetGen (an isotropic boundary-constrained Delaunay-based method [61]) is used to tetrahedralize the remainder of the blood region. Boundary layers in meshes are useful when calculating metrics such as pressure or wall shear stress (WSS) caused by blood flow [52, 16]. Although the method generated high-quality hybrid meshes, the Delaunay-based isotropic elements in fluid flow regions (such as in the coronary arteries) were identified as a drawback, with anisotropy suggested as an improvement. Additionally, Delaunay tetrahedralization limits the quality of the tetrahedral core for complex biological geometries [16] and can fail to terminate for certain cases [27]. Another method was developed to model cerebral aneurysms for fluid-structure interaction simulations [66]. The surface mesh of a healthy blood vessel is deformed onto an aneurysm model, simulating the formation of the aneurysm. While the simulation results demonstrated the benefit of utilizing anisotropic meshes rather than isotropic meshes, the mesh generation method is sequential, as with all the aforementioned adaptive anisotropic methods.
The proposed approach combines multiple software tools into a single pipeline to provide the following: (1) image-to-mesh conversion which satisfies the quality, fidelity, and smoothness ($C^{0}$) requirements, (2) the generation of a boundary layer grid over the high fidelity surface, (3) a parallel adaptive anisotropic meshing procedure which satisfies the real-time requirement, and (4) robustness, which is satisfied by the pipeline’s ability to process segmented images and CAD models (such as that shown in fig. 2 [64]). To the best of our knowledge, there is no single method that provides all of the above.
3 Background.
3.1 Image-to-mesh Conversion.
Image-to-mesh conversion is handled by a software called CBC3D [27]. CBC3D converts a segmented, multi-labeled image into an isotropic tetrahedral (or mixed element) mesh. The segmented image is discretized using a regular Body-Centered Cubic lattice. The user has the option of adjusting the distance between the vertices of the lattice. Any resulting tetrahedra located outside the object are discarded. Adaptive refinement commences over these high quality elements using red (regular) - green (irregular) subdivisions. A Euclidean Distance Transform is used to determine if an element should be subdivided. CBC3D allows the user to specify an input fidelity parameter which influences how fine the mesh will be near the boundaries. A label-specific fidelity can also be specified to allow for localized refinement in materials of interest (reducing overall element count). Additionally, a mixed element mesh can be created where tetrahedra are merged into hexahedra. The topology of generated surfaces or interfaces between labels is preserved by only merging within homogenous regions where the lattice is uniform. While useful for further reducing element count, the proposed approach only utilizes tetrahedral meshes generated by CBC3D given that CDT3D (the adaptive anisotropic method) only processes tetrahedral background meshes. As seen in our evaluation (see section 5), CDT3D meets the real-time requirement when processing tetrahedral background meshes. The adaptive refinement approach utilized by CBC3D guarantees an output mesh with high quality elements [27]. CBC3D finally deforms the mesh surfaces to their corresponding image boundaries. The final mesh has a smoothness of $C^{0}$ while maintaining the same high quality. Consequently, CBC3D was selected to be the image-to-mesh conversion tool within the pipeline of the proposed approach because it: (1) generates meshes of high quality and fidelity, (2) provides smoothness, and (3) was previously used to discretize patient-specific medical imaging data and STLs of CAD-based stent geometries for use within CFD simulations [40, 26, 25].
3.2 Boundary Layer Generation.
The generation of boundary layers over the high fidelity CBC3D mesh surface is carried out by a software known as AFLR. AFLR is an unstructured mesh generation method used extensively in industry. While it is capable of both isotropic [49] and anisotropic mesh generation [47], it is sequential. Consequently, its use within the proposed approach is only focused on its boundary layer generation capability [48]. AFLR generates boundary layers using a spacing that is normal to the boundary and that grows geometrically. As new points are generated, quality is maintained by checking their distribution (i.e., distance from one another according to a local element length scale). New points are only accepted and inserted into the grid if they produce elements that satisfy dihedral angle and element aspect ratio thresholds. The user has numerous options for controlling the spacing distribution and may also specify the number of layers generated.
3.3 Adaptive Anisotropic Mesh Generation.
The final adaptive anisotropic mesh is generated using a software tool called CDT3D. CDT3D is a multicore cc-NUMA-based mesh generation method that exploits fine-grain parallelism at the cavity level using data decomposition. It is capable of both isotropic [28] and adaptive anisotropic mesh generation [64]. A speculative execution model is implemented, which serves to exploit parallelism “everywhere possible” throughout the mesh generation procedure. Atomic lock instructions are utilized by several of CDT3D’s meshing operations so that different data can be modified concurrently while guaranteeing correctness (i.e., conformity). A lock attempts to acquire the necessary dependencies for a corresponding operation. If unable to do so, any acquired resources are released and the operation is applied on a different set of data. This is repeated until the mesh satisfies spacing and qualitative criteria. There are a number of parallel mesh operations in CDT3D, including: edge collapse, point creation/insertion (which by default uses an advancing front technique for isotropic mesh generation and a centroid-based technique for anisotropic adaptation), CAD projection, local reconnection (edge/face swapping), and point smoothing. See [28] and [64] for more information on their design and implementation.
CDT3D accepts an analytic or discrete metric field as input when performing metric-based anisotropic mesh adaptation. CAD geometries can also be processed and used to perform adaptation (thus allowing the proposed approach to handle both medical I2M conversion and the processing of CAD models for CFD simulations involving stents). CDT3D has been quantitatively and qualitatively evaluated with several benchmarks based on aerospace cases [65, 64]. Not only was it shown to be stable in terms of metric conformity (element shape size and edge length), it also showcased excellent performance when utilizing up to 40 cores on a single multicore node. When coupled with a solver in a CFD simulation pipeline, it captured features of the underlying simulation and occupied only a small fraction of the pipeline’s runtime [64] (a visual example is shown in fig. 2).
4 Method.
The proposed approach involves several steps: (1) utilize CBC3D to convert a segmented image into a high quality and high fidelity background mesh (for anisotropic mesh adaptation), (2) construct an anisotropic metric tensor field, (3) generate a boundary layer grid outside the CBC3D mesh surface using AFLR, and (4) utilize CDT3D to generate an anisotropic mesh volume using the metric tensor field. A visual overview of this pipeline of tools can be seen in fig. 3.
Figure 3: Presented is the pipeline of software tools used to satisfy the medical image-to-mesh conversion requirements for use within a CFD vascular flow simulation. It begins with image-to-mesh conversion (handled by CBC3D). Extracting the medial axis with VMTK [9, 8] is optional, as the method in [21] is only used to build an anisotropic metric tensor field for the purpose of testing this pipeline’s capability to meet the aforementioned requirements (described in section 5.1). A boundary layer grid is generated over the CBC3D mesh using AFLR while the adaptive anisotropic mesh volume is generated by CDT3D (using the CBC3D mesh as a background mesh). The boundary layer grid and anisotropic mesh are merged into one final mesh.
The proposed approach is not bound to any particular method for constructing the metric tensor field. As stated previously, the pipeline is evaluated based on the aforementioned requirements to gauge its suitability for CFD vascular flow simulations. Consequently, our method is tested using two different techniques for constructing a metric tensor field, where each is used with a different anuerysm case. These techniques are described in our experimental setup, section 5.1. The anisotropic metric tensor field and CBC3D background volume mesh (converted from the segmented image) are provided as input to CDT3D, which generates an adaptive anisotropic volume mesh. In this study, surface adaptation is turned off to ensure that the high fidelity surface generated by CBC3D remains intact. The surface is extracted from the CBC3D mesh and is given as input to AFLR to generate a boundary layer grid. AFLR is set to only generate a boundary layer grid, designating the input as an open domain so that only boundary layer elements are generated on the outside of the surface. Surface refinement is also turned off for AFLR (again, for this particular study) to ensure that the high fidelity CBC3D surface remains intact and so that both the boundary layer grid and CDT3D mesh are conforming. The two are merged to create one final output mesh.
5 Results.
5.1 Experimental Setup.
The first aneurysm case is based on a rotational angiography scan of a carotid cavernous aneurysm with image spacing of 1.00 x 1.00 x 1.00 mm3 and an image size of 512 x 512 x 508 voxels3. The CBC3D mesh generated from it can be seen in LABEL:fig:artery1, in addition to its approximate medial axis. The second case is a surface mesh of a middle cerebral artery bifurcation aneurysm given as a Piecewise Linear Complex (PLC) in a .stl format (seen in LABEL:fig:artery2_plc_2). The full pipeline of the proposed approach is tested with the first case while each method attempts to perform mesh generation/adaptation using only the surface mesh for the second case. All codes were compiled using GNU GCC 7.5.0. Data were collected on Old Dominion University’s Wahab cluster using a dual socket node that features two Intel® Xeon® Gold 6148 CPUs @ 2.40 GHz (20 slots) and 384GB of memory. For a performance evaluation, each run in the following experiments was executed five times and the results were averaged.
We evaluate the adaptive anisotropic mesh generation method with metric tensor fields that are constructed using different techniques. For the first case, an anisotropic metric tensor field is constructed based on a method also utilized for non-rigid registration applications intended for image-guided neurosurgery [21]. This method focuses on capturing registration points within an adaptive mesh. A metric tensor field is built by evaluating the k-closest registration points to each mesh vertex using a k-nn search from the VTK library [6]. The Khachiyan algorithm is then used to construct a minimum volume bounding ellipsoid which encompasses the given point set. This ellipsoid has a natural mapping to a 3 x 3 positive definite matrix that can be used as a metric to guide adaptivity. See [21] for more details. To exploit parallelism, an OpenMP construct is utilized when processing the mesh points. For our first test case, we utilize this method using the points of an approximate medial axis of the generated CBC3D mesh (in place of registration points). The approximate medial axis (also called a centerline) is obtained for our first test case using VMTK [9, 8] and can be seen in LABEL:fig:artery1.
LABEL:fig:artery2_images shows the vascular structure imaged by rotational angiography of the right carotid artery for the second aneurysm case, in addition to the surface mesh (utilized in this study) obtained from a segmentation of the region of interest. LABEL:fig:artery2_streamlines shows streamline data from a CFD simulation for this case (executed outside of this study [43]). The velocity at each point along these streamlines is used to construct a metric tensor field to guide adaptivity for anisotropic mesh generation. A velocity gradient tensor is calculated at each point which can be mapped to a 3x3 positive definite matrix. The magnitude (eigenvalue) of each velocity gradient tensor is used to apply a stretching factor to the positive definite matrix which ensures that elements are stretched in the dominant direction of the velocity. An isotropic volume mesh is first generated using the surface mesh of the second case. A k-nn search is utilized to find the 3 closest velocity points to each mesh volume vertex (which are then used to calculate the velocity gradient tensor for each mesh point). OpenMP is again utilized to parallelize the processing of the mesh points. The isotropic volume mesh is processed as a background mesh with the velocity-based metric tensor field to generate anisotropic meshes. Both the anisotropic and isotropic meshes generated by CDT3D for the second case are evaluated.
5.2 Quantitative and Qualitative Analysis
The results obtained from CDT3D are also compared to TetGen [61, 5] (replacing the adaptive anisotropic method with this adaptive isotropic method in the pipeline). TetGen is a sequential, Delaunay-based method that generates isotropic tetrahedral meshes. In addition to an anisotropic metric field, an isotropic sizing function is also constructed (based on the aforementioned approach in [21]) specifically for TetGen. We compare results between meshes generated by TetGen when it utilizes this sizing function and when it does not (tetrahedralization of the CBC3D surface). Additionally, we compare the TetGen isotropic meshes to isotropic meshes generated by CDT3D (again simply processing the same surface as a PLC - .stl).
Table 1 shows the parameters used for each mesh generation method in the pipeline (except CBC3D). The first aneurysm case was also used to evaluate CBC3D in [27] (as it was compared to other isotropic I2M conversion methods from industry and academia). The same parameters utilized for CBC3D in that study are utilized in this evaluation. A wide range of values were tested for the value of k used in the k-nn search when building the sizing function and metric tensor fields for the first case. The optimal parameters were determined based on the best quality mesh that was generated. If certain parameters are not listed in table 1, then their default value was utilized. The optimal value for k with the isotropic runs of TetGen is 30. The optimal value for k in the metric-adapted anisotropic runs of CDT3D is 40. The CDT3D parameters nqual and nsmth were both increased to 10 while cdfn was changed to 0.8 to improve the final anisotropic mesh quality for the second aneurysm case (with no surface adaptation by CDT3D, explained later).
Table 1: Input parameters for each mesh generation method utilized in the pipeline (after CBC3D). Val. is an abbreviation for value.
Method Parameter/Val. Description
TetGen (iso-sizing) r reconstruct mesh
m use sizing
q optimize quality
R reduce elements
O/9 optimize mesh
Y freeze surface
TetGen (iso) p mesh plc
q improve quality
R reduce elements
O/9 optimize mesh
Y freeze surface
AFLR (bl) mbl/1 boundary layer
blds/1.2 initial bl spacing
(for case 1)
blds/0.12 initial bl spacing
(for case 2)
fints/0, freeze surface
open only bl
CDT3D (iso) sref/0 freeze surface
nthreads/* parallel threads
CDT3D (aniso) met/1 use metric
meval/1 metric-vertices
sref/0 freeze surface
reconstruct/1 reconstruct mesh
mpp/5 point creation
method - centroid
cdff/1.2 edge length mult.
cdfn/0.6 nearby node dist.
nqual/3 optimize quality
nsmth/5 smooth points
nthreads/* parallel threads
CBC3D takes approximately 61.59 seconds (1 min) to convert the segmented image of the first aneurysm into a mesh of approximately 272K tetrahedra and 63K points. LABEL:fig:Mesh_Cuts_artery1 shows cross sections of the volume meshes generated by each method, merged with the conforming AFLR boundary layer mesh. LABEL:fig:Mesh_Cuts_artery2 shows cross sections of the volume meshes generated for the second aneurysm case. Table 2 shows the number of elements generated by each method when processing the CBC3D mesh of aneurysm 1 and the surface mesh of aneurysm 2. The TetGen meshes contain fewer elements compared to the CDT3D meshes in both cases. Table 3 and table 4 show the runtime of each method for aneurysm cases 1 and 2, respectively. They also show the mesh generation/adaptation rates of each method (measured in elements generated per second). TetGen exhbits good performance when generating its small-size meshes. Although it is not listed in table 3, it should be noted that CDT3D exhibits good scalability, as its runtime is reduced to approximately 1 minute and its anisotropic adaptation rate becomes 27K when utilizing 40 cores. When utilizing 10 cores, CDT3D’s mesh generation/adaptation rate exceeds that of TetGen’s. For the second case, CDT3D completes anisotropic adaptation in about 49 seconds when utilizing 40 cores.
Table 2: Number of tetrahedra and points generated by each method when processing the CBC3D mesh converted from the segmented image of the first aneurysm and the surface mesh of the second aneurysm. ‘Tets’ is short for tetrahedra. K means thousand and M means million.
Methods Aneurysm 1 Aneurysm 2
TetGen (iso-sizing) Tets 120K -
Points 37K -
TetGen (iso) Tets 118K 12K
Points 36K 3K
AFLR (bl) Tets 805K 105K
Points 144K 18K
CDT3D (iso) Tets 631K 31K
Points 122K 7K
CDT3D (aniso) Tets 1.59M 80K
Points 284K 15K
Table 3: The time spent (approximately in minutes) and the number of elements generated per second by each method (in parentheses) for small-size meshes is shown when processing the CBC3D mesh converted from the segmented image of the first aneurysm. Instead of elements per second, the values in parentheses for Build_Sizing/Metric represent the CBC3D volume mesh points processed per second. K means thousand.
CPU Cores
Methods 1 10 20
Build_Sizing/Metric 191 (5K) 19 (55K) 9 (117K)
TetGen (iso-sizing) 0.1 (20K) - -
TetGen (iso) 0.06 (32K) - -
AFLR (bl) 0.2 (67K) - -
CDT3D (iso) 0.6 (17K) 0.3 (42K) 0.1 (78K)
CDT3D (aniso) 20 (1K) 3 (8K) 2 (14K)
Table 4: The time spent (approximately in seconds) and the number of elements generated per second by each method (in parentheses) for small-size meshes is shown when processing the surface mesh of the second aneurysm. K means thousand.
CPU Cores
Methods 1 10 20
TetGen (iso) 1.54 (8K) - -
AFLR (bl) 1.8 (58K) - -
CDT3D (iso) 2.2 (14K) 1.73 (18K) 1.24 (25K)
CDT3D (aniso) 910 (87) 105 (761) 69 (1K)
Mesh fidelity is evaluated using a two-sided Hausdorff Distance (HD) metric (based on an open source implementation [23]): $\text{HD}=\max\{\text{HD}_{\text{I}\rightarrow\text{M}},\text{HD}_{\text{M}\rightarrow\text{I}}\}$, where $\text{HD}_{\text{I}\rightarrow\text{M}}$ is the value of the metric from the image to the mesh, and $\text{HD}_{\text{M}\rightarrow\text{I}}$ is the value of the metric from the mesh to the image. A low HD error indicates a high fidelity. Hausdorff Distance is computed between two point sets. The first point set contains the vertices located on the mesh surface and the second point set contains the voxels located on the boundaries of the segmented material. For the first aneurysm case (in LABEL:fig:artery1), $\text{HD}_{\text{I}\rightarrow\text{M}}$ is 4.51 while $\text{HD}_{\text{M}\rightarrow\text{I}}$ is 1.74. The HD is therefore 4.51. The fidelity of the mesh generated by CBC3D was also evaluated for the first aneurysm case in [27]. CBC3D was compared to several highly used I2M methods, exhibiting good fidelity while maintaining better smoothness compared to the other methods’ meshes. Surface refinement/adaptation is turned off for all mesh generation methods tested in this pipeline (except CBC3D). This ensures that the high fidelity and smooth surface generated by CBC3D remains intact.
For the isotropic meshes, qualitative results are examined with respect to element dihedral angles. LABEL:fig:artery1_dihedrals compares the quality of the isotropic meshes generated by each method for case 1 and LABEL:fig:artery2_dihedrals compares those generated for case 2. Table 5 shows the smallest and largest dihedral angles found in the isotropic meshes generated by each method. The CDT3D meshes contain better quality compared to the TetGen meshes. This is particularly seen in (b) and (c) of LABEL:fig:artery1_dihedrals and LABEL:fig:artery2_dihedrals, which show that the TetGen meshes contain more elements with small and large angles. For the adapted anisotropic meshes, qualitative results are examined with respect to their metric conformity. The goal of metric conformity is to create a unit grid, where edges are unit-length and elements are unit-volume with respect to the target metric. For calculating edge length, we adopted the same definition that appears in [7]. For two vertices a and b, an edge length in the metric $L_{e}$ can be evaluated using:
(5.1) $\begin{split}&L_{e}=\begin{cases}\frac{L_{a}-L_{b}}{log(L_{a}/L_{b})}&|L_{a}-L_{b}|>0.001\\ \frac{L_{a}+L_{b}}{2}&otherwise\end{cases}\\ &L_{a}=(v_{e}^{T}M_{a}v_{e})^{\frac{1}{2}},L_{b}=(v_{e}^{T}M_{b}v_{e})^{\frac{1}{2}}\end{split}$
and an element mean ratio shape measure can be approximated in the discrete metric as:
(5.2) $Q_{k}=\frac{36}{3^{1/3}}\frac{\left(|k|\sqrt{det(M_{mean})}\right)^{\frac{2}{3}}}{\sum_{e\epsilon L}v_{e}^{T}M_{mean}v_{e}},$
where v is a vertex of element k and $M_{mean}$ is the interpolated metric tensor evaluated at the centroid of element k. Given that optimal edges should be unit-length, edges with length above or below one are considered to be sub-optimal. The measure for mean ratio is bounded between zero and one since it is normalized by the volume of an equilateral element. One is the optimal quality for an element’s mean ratio shape. LABEL:fig:arteries_CDT3D_MR_EL shows that CDT3D maintains good quality when generating an anisotropic mesh for the first case. Although there are 3 elements with a mean ratio measure less than 0.1 (the minimum shape measure is 0.016), the majority of elements have a high mean ratio and edge lengths close to 1. With regards to the second case, the anisotropic mesh generated by CDT3D exhibits poor quality when the surface of the input mesh is not adapted (containing about 1K elements with a mean ratio less than 0.1). However, when CDT3D is permitted to adapt the surface, the final mesh quality is much better (about 20 elements have a mean ratio less than 0.1, with the minimum shape measure as 0.042). When adapting the surface, the same parameters as those in table 1 are utilized with surface adaptation turned on. It should be noted that CDT3D converges and completes anisotropic adaptation in less time when surface adaptation is permitted (generating 90K tetrahedra in about 22 seconds when utilizing 20 cores) as opposed to its performance when restricting the method to preserve the input surface. The quality of the CDT3D mesh generated for the first case shows that CBC3D indeed generates a mesh of good quality when converting the first case image, given that surface adaptation is turned off for CDT3D and it still generates a final volume of good quality. In order to provide a volume that conforms to the boundary layer grid generated by AFLR, surface adaptation is turned off and CDT3D is therefore limited by the quality of the input surface mesh for the second aneurysm case. It should be noted that surface adaptation was attempted first, where the CDT3D-adapted surface was provided as input to AFLR. However, AFLR failed to process the CDT3D-adapted surface and generate a boundary layer grid. We also attempted surface refinement and volume generation using AFLR, to then subsequently use as a background mesh for CDT3D’s anisotropic adaptation (which could potentially be merged with a boundary layer grid generated over the AFLR-refined surface). AFLR however fails to refine the input PLC surface of the second case. Nevertheless, the majority of elements in both anisotropic meshes generated by CDT3D for the second aneurysm case (with and without surface adaptation) have a high mean ratio and edge lengths close to 1.
Table 5: Smallest and largest dihedral angles in isotropic meshes generated by TetGen and CDT3D for both aneurysm cases.
Methods Angle Aneurysm 1 Aneurysm 2
TetGen (iso-sizing) Smallest 1.67 -
Largest 176.73 -
TetGen (iso) Smallest 3.99 2.71
Largest 171.66 171.3
CDT3D (iso) Smallest 3.99 3.54
Largest 165.25 172.93
6 Discussion and Future Work.
Although it doesn’t generate anisotropic meshes, TetGen exhibits good quality while generating much fewer elements compared to CDT3D (for the small test cases of our evaluation). Additionally, TetGen meets the real-time requirement for small-size problems. CDT3D generates meshes with better quality but at the cost of generating more elements. Depending on the accuracy needed for CFD simulations, this tradeoff may be acceptable, particularly when combined with CDT3D’s anisotropy (seen in LABEL:fig:Mesh_Cuts_artery1 and LABEL:fig:Mesh_Cuts_artery2). Both the Build_Sizing/Metric and CDT3D methods show good scalability on a single multicore node, particularly when generating an anisotropic mesh. Due to the fact that CDT3D is designed to handle much larger problems (such as that in fig. 2 and LABEL:fig:artery2_images), its performance suffers from the overhead of parallelizing adaptation for such small meshes. Nevertheless, it also meets the real-time requirement when utilizing a full multicore node (for anisotropic mesh adaptation) and when utilizing a single CPU core for isotropic mesh generation. CDT3D exhibits much faster performance when generating isotropic meshes from the PLCs of the first and second cases.
The parallelism offered by CDT3D would be well suited for processing a geometry generated from an image of a large vascular structure (such as in LABEL:fig:artery2_images). This would likely require many more elements compared to the small test cases of this study. Consequently, additional memory would also be required, potentially necessitating the need for a distributed memory approach (such as [35]) if a sequential or multicore cc-NUMA-based method cannot generate and store such a large mesh within the memory of a single machine (millions of elements, or potentially hundreds of millions of elements which in the future may be needed for aneurysm cases requiring even finer mesh resolution, e.g., microvascular aneurysms). Table 3 shows the elements generated per second up to 20 cores by CDT3D for the first aneurysm case. As mentioned previously, the adaptation rate becomes 27K when utilizing 40 cores. For comparison, TetGen generates 20K isotropic elements per second (when using the sizing function for adaptivity). Consider an example where a vascular structure requires a mesh of 10 million elements in order to obtain the needed accuracy in a CFD simulation. Based on the above rates, TetGen would take approximately 8 minutes (or 83 minutes in the case of a 100 million element mesh) to generate an adaptive isotropic mesh while CDT3D would take approximately 6 minutes (or 61 minutes for 100 million elements) to generate an adaptive anisotropic mesh using 40 cores. Furthermore, CDT3D utilizes a similar speculative execution model (described in section 3.3) that was utilized within the I2M method in [31]. With an additional hierarchical work-stealing optimization for load balancing, this model for parallelization was tested and shown to maintain a speedup of approximately 123 when executed on a distributed shared memory architecture using 144 cores (seen in table 4 of [31]). If CDT3D were to receive a similar optimization and be executed on a similar architecture, it could potentially achieve an adaptation rate of 98K elements per second (based on the rates seen thus far in table 3), generating approximately 10 million anisotropic elements in 1.7 minutes (100 million elements in 17 minutes). CDT3D’s adaptation rate, coupled with its scalable performance and better quality, may prove beneficial when anisotropy is needed to provide better simulation results compared to isotropy (such as in [66, 16]). Future work is required to optimize and test CDT3D on a distributed shared memory architecture in order to gauge its potential and verify this performance estimation.
Although the methods utilized in the proposed approach have been shown to satisfy the I2M requirements using the small test cases in our evaluation, additional experimentation (and fine-tuning of each method’s parameters) is required to process larger test cases (such as the complete vascular structure in LABEL:fig:artery2_original_image_1). The mesh deformation scheme employed by our I2M approach produces meshes with a smoothness of $C^{0}$. This smoothness is sufficient for aneurysm geometries such as those in our evaluation [27]; however, future work is required to improve the smoothness of geometries reconstructed from micro-CT images (i.e., some stent geometries). Our approach must be updated to provide a smoothness of $C^{1}$. Given CDT3D’s ability to process CAD models (such as in fig. 2), this approach remains to be tested within a CFD vascular flow simulation that includes CAD models of stents (testing potential avenues of treatment for other aneurysm cases).
Eventually, the tools utilized in the pipeline of the proposed approach should be integrated into a single software. It should be noted that conversions are required between the input and output files of the methods utilized. For example, several of the methods use the VTK library [6]. Because they were each developed at different time periods, one method produces output utilizing a newer version of the library while another (which utilizes an older version) needs to read this as input. Although the conversion times for the small test cases in our evaluation are negligible, this discrepancy between the methods will likely negatively impact performance when the pipeline is tested on large cases. Integrating these methods into a single software will not only eliminate this problem, but will remove the complexity of compiling all codes separately and running them one after the other in the pipeline. Rather than separating image-to-mesh surface adaptation from anisotropic volume generation (and freezing the surface during this step), it would be more effective to perform both simultaneously when converting the segmented image (similar to the isotropic approach in [31]). If integrated with a boundary layer generation method, this can help prevent the inhibition of mesh quality improvement (such as that shown for the second aneurysm case in LABEL:fig:arteries_CDT3D_MR_EL). Finally, the sequential methods must be parallelized (or replaced with other parallel methods) in order to maximize potential performance, especially for large cases.
7 Conclusion.
Presented is a path towards a fast and robust adaptive anisotropic mesh generation method that is designed to help streamline the discretization of complex vascular geometries and CAD-based stent models within the CFD modeling process. The proposed approach combines multiple software tools into a single pipeline to provide the following: (1) image-to-mesh conversion which satisfies quality, fidelity, and smoothness requirements, (2) the generation of a boundary layer grid over the high fidelity surface, (3) a parallel adaptive anisotropic meshing procedure which satisfies the real-time requirement, and (4) robustness, which is satisfied by the pipeline’s ability to process segmented images and CAD models. This pipeline is tested on a segmented image of a carotid cavernous aneurysm. Additionally, the parallel meshing method is also tested using a surface mesh of a middle cerebral artery bifurcation aneurysm. Both aneurysm cases were obtained from the Cerebrovascular Center for Research at Stony Brook University under Institutional Review Board approval. The pipeline is shown to satisfy all the aforementioned requirements. More specifically, the parallel mesh generation method successfully produces high quality isotropic and anisotropic meshes for the aneurysm cases in real-time. Having shown the feasibility of this approach with regards to these requirements, the next steps are to fully parallelize the remaining components of the pipeline to maximize potential performance and to test its integration within a CFD vascular flow simulation. Just as the parallel anisotropic adaptation procedure was tested within aerospace CFD simulations using CAD models (such as that shown in fig. 2 [64]), the method is expected to provide accurate results for CFD vascular flow simulations in real-time when executed on multicore cc-NUMA architectures.
Acknowledgements
This research was sponsored by the Richard T. Cheng Endowment at Old Dominion University, SURA grant No. C2024-FEMT-011-01, and the National Institute of General Medical Sciences of the National Institutes of Health under Award Number 1T32GM140911-03. This study was also partially supported by a pilot award from the Long Island Network for Clinical and Translational Science. The content is solely the authors’ responsibility and does not necessarily represent the official views of the National Institutes of Health. Additionally, this work was supported by the research computing clusters at Old Dominion University. The authors would like to thank Dr. Christos Tsolakis for helping to assemble the tools utilized in the adaptive pipeline of our approach. We’d also like to thank Dr. David Marcum for sharing the AFLR software utilized in this study, and Evangelia Frastali for helping to determine the best parameters for each method tested in our evaluation.